OpenStudy (anonymous):

14a + 10a = 72 please show all work

OpenStudy (anonymous):

24a = 72 a = 3

OpenStudy (anonymous):

14a + 10a = 72 24a=72 24a/24=72/24 a=3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)